# FerretNet: Efficient Synthetic Image Detection via Local Pixel Dependencies

### v1

@@ Abstract
The increasing realism of synthetic images generated by advanced models such as VAEs, GANs, and LDMs poses significant challenges for synthetic image detection. To address this issue, we explore two artifact types introduced during the generation process: (1) latent distribution deviations and (2) decoding-induced smoothing effects, which manifest as inconsistencies in local textures, edges, and color transitions. Leveraging local pixel dependencies (LPD) properties rooted in Markov Random Fields, we reconstruct synthetic images using neighboring pixel information to expose disruptions in texture continuity and edge coherence. Building upon LPD, we propose FerretNet, a lightweight neural network with only 1.1M parameters that delivers efficient and robust synthetic image detection. Extensive experiments demonstrate that FerretNet, trained exclusively on the 4-class ProGAN dataset, achieves an average accuracy of 97.1% on an open-world benchmark comprising across 22 generative models, surpassing state-of-the-art methods by 10.6%.
@@ 1 Introduction
Many existing detection approaches rely heavily on model-specific features, which limits their generalization ability to unseen generative architectures. For example, Durall et al. [7] observed characteristic frequency artifacts in GAN-generated images. Although frequency-domain techniques [50, 15, 16] have demonstrated strong performance under known conditions, they often struggle to generalize across different models. DIRE [51] introduced a diffusion-based detection framework that distinguishes synthetic images by reconstructing them through a diffusion model, a capability that fails with real images. However, this method performs poorly when applied to GAN-generated content.
@@
Based on these insights and grounded in the theory of Markov Random Fields, we introduce a pixel-level artifact representation that captures local pixel dependencies (LPD) through median-based reconstruction. We further propose FerretNet, a lightweight detector designed with depthwise separable and dilated convolutions to enhance both performance and efficiency.
@@
  • We construct a large-scale benchmark dataset, Synthetic-Pop, comprising 30,000 synthetic images generated by six different models and 30,000 real images from COCO [23] and LAION-Aesthetics V2 (6.5+) [43], totaling 60,000 images.
@@ 2.1 Pixel-based Synthetic Image Detection
Wang et al. [50] trained a classifier on images generated by a single model to detect fake images across various architectures and datasets, addressing cross-model generalization via data augmentation and diverse training samples. Shi et al. [46] proposed a difference-guided reconstruction learning framework that exploits discrepancies between real and synthetic images to enhance detection accuracy. Ojha et al. [29] tackled the generalization problem to unseen generative models by leveraging a feature space not explicitly trained for real/fake discrimination, employing nearest-neighbor and linear probing strategies. He et al. [12] introduced a super-resolution-based re-synthesis technique to reconstruct test images and extract residual or layered artifact features, thereby reducing reliance on frequency artifacts. Tan et al. [49] proposed NPR, a method that revisits the upsampling process in generative CNNs by modeling Neighbor Pixel Relations, aiming to improve generalization in deepfake detection. Liu et al. [24] designed a robust detection framework based on multi-view image completion, which simulates real image distributions and captures frequency-independent features. FatFormer [25] presented a forgery-aware adaptive transformer incorporating forgery-specific adapters and language-guided alignment modules to better adapt pre-trained models for synthetic image detection.
@@ 2.2 Frequency-based Synthetic Image Detection
F3Net [35] introduced a dual-branch architecture that captures frequency-aware clues for detecting subtle forgery traces, particularly in low-quality and facial imagery. FrePGAN [16] developed a frequency-level perturbation GAN framework, where a generator-discriminator pair is used to iteratively improve classifier robustness against unseen categories and generative models. Tan et al. [47] exploited pre-trained CNN gradients to generate generalizable representations of GAN-specific artifacts. BiHPF [15] amplified frequency-level artifacts via a high-pass filtering approach, achieving improved robustness across diverse image categories, color manipulations, and generative models. FreqNet [48] introduced high-frequency representations and frequency-specific convolution layers to enhance detection by focusing on localized high-frequency components, addressing overfitting and poor generalization seen in prior methods.
@@ 4.1 Local Median-based Feature Extraction
We propose a synthetic image detection method based on local statistical dependencies. The core idea is to identify generation artifacts by quantifying the deviation of each pixel from the median of its surrounding neighborhood. The full computational procedure is outlined in Algorithm 1.
@@ 5.1.1 Training Dataset
To ensure a consistent evaluation baseline, we follow the protocols established in [15, 29, 25, 49], utilizing four semantic classes (car, cat, chair, horse) from the ForenSynths dataset [50]. Each class contains 18,000 synthetic images generated by ProGAN [18], paired with an equal number of real images from the LSUN dataset [55]. In all experiments conducted in this study, only the aforementioned 4-class data (i.e., the 4-class ProGAN dataset) was used for training; no additional data sources were included.
@@ 5.3 Main Results
Dataset Wang [50] F3Net [35] LGrad [47] Ojha [29] FreqNet [48] NPR [49] FatFormer [25] FerretNet
Dall-E 51.8/61.3 71.6/79.9 88.5/97.3 89.5/96.8 97.3/99.7 90.9/98.1 98.8/99.8 91.4/98.2
Guided 54.9/66.6 69.2/70.8 86.6/100.0 75.7/85.1 67.2/75.4 74.0/78.1 76.1/92.0 92.1/98.6
PNDM 50.8/90.3 72.8/99.5 69.8/98.5 75.3/92.5 85.2/99.9 97.5/100.0 99.3/100.0 96.9/100.0
VQ-Diffusion 50.0/71.0 100.0/100.0 96.3/100.0 83.5/97.7 100.0/100.0 100.0/100.0 100.0/100.0 99.9/100.0
Glide-50-27 54.2/76.0 88.5/95.4 90.7/95.1 91.1/97.4 86.6/95.8 97.5/99.5 94.7/99.4 97.2/99.7
Glide-100-10 53.3/72.9 88.3/95.4 89.4/94.9 90.1/97.0 87.8/96.0 97.8/99.5 94.2/99.2 97.9/99.9
Glide-100-27 53.0/71.3 87.0/94.5 87.4/93.2 90.7/97.2 84.4/95.6 97.4/99.5 94.4/99.1 97.3/99.7
LDM-100 51.9/63.7 74.1/84.0 94.8/99.2 90.5/97.0 97.8/99.9 98.0/99.6 98.7/99.9 98.8/100.0
LDM-200 52.0/64.5 73.4/83.3 94.2/99.1 90.2/97.1 97.4/99.9 98.2/99.6 98.6/99.8 98.8/100.0
LDM-200-CFG 51.6/63.1 80.7/89.1 95.9/99.2 77.3/88.6 97.3/99.9 98.0/99.5 94.9/99.1 98.5/99.9
Mean 52.4/70.1 80.6/89.2 89.4/97.7 85.4/94.6 90.1/96.2 94.9/97.3 95.0/98.8 96.9/99.6
@@
Methods Parameters Throughput ACC / AP on the Test set
(M) $\downarrow$ (Img / s) $\uparrow$ ForenSynths Diffusion-6-cls Synthetic-Pop Synthetic-Aesthetic Mean
FreqNet [48] 1.9 200.2 91.5 / 98.5 90.1 / 96.2 65.0 / 71.4 70.1 / 81.2 79.2 / 86.8
NPR [49] 1.4 720.9 92.5 / 96.1 94.8 / 97.7 77.4 / 81.2 81.4 / 82.4 86.5 / 89.4
FatFormer [25] 577.3 88.6 98.4 / 99.7 95.0 / 98.8 70.5 / 74.8 80.4 / 90.6 86.1 / 91.0
FerretNet 1.1 772.1 95.9 / 99.3 96.9 / 99.6 98.3 / 99.8 97.3 / 99.6 97.1 / 99.6
Real
CAM
Fake
CAM
Figure 4: Grad-CAM Visualizations: The fake images are synthesized using RealVisXL-4.0, SDXL-Turbo, SD-3.5-Medium, StyleGAN, and BigGAN. FerretNet show stronger activation responses in synthetic images, while showing no significant responses in real ones.
@@
Next, on diffusion-generated images (Table 2), FerretNet attains an ACC of 96.9% and an AP of 99.6%, outperforming FatFormer [25] by 1.9 and 0.8 percentage points (pp), respectively. Other lightweight models such as NPR [49] and FreqNet [48] perform less favorably, with ACC scores falling below 95.0%.
We further evaluate performance on high-quality synthetic images using the Synthetic-Pop test set (Table 3). Existing methods experience noticeable degradation; for example, NPR [49]—the strongest among them—achieves only 77.4% ACC and 81.2% AP. In contrast, FerretNet maintains 98.3% ACC and 99.8% AP, highlighting its robustness and reliability on visually realistic forgeries.
To evaluate real-world applicability, we tested FerretNet on the Synthetic-Aesthetic set for both detection performance and efficiency. As shown in Table 4, FerretNet achieves 97.3% ACC and 99.6% AP with 1.1M parameters and 772.1 FPS on an RTX 4090. While comparisons are made on this set, its average performance across all four benchmarks reaches 97.1% ACC and 99.6% AP. Notably, it outperforms FatFormer [25] by 11.0 and 8.6 pp in ACC and AP, respectively, while using only 0.2% of its parameters.
Finally, Grad-CAM [44] visualizations in Figure 4 show that FerretNet focuses on pixel-level artifacts rather than high-level semantics, contributing to its strong generalization to unseen generators.
@@ 5.4 Ablation Study
Unless specified, all ablation results report the average ACC and AP across four datasets: ForenSynths, Diffusion-6-cls, Synthetic-Pop, and Synthetic-Aesthetic.
@@ 6 Conclusion
This work presents a universal artifact representation framework and introduces FerretNet, a lightweight yet effective neural network for synthetic image detection. FerretNet achieves a remarkable 99.8% reduction in parameters compared to state-of-the-art methods FatFormer [25], while maintaining exceptional detection accuracy, reaching 97.1% on images generated by 22 different generative models. It demonstrates strong generalization capabilities and computational efficiency, outperforming existing approaches on high-quality synthetic datasets. Our contributions include a novel artifact representation approach and the introduction of the Synthetic-Pop dataset.

### v2

@@ Abstract
The increasing realism of synthetic images generated by advanced models such as VAEs, GANs, and LDMs poses significant challenges for synthetic image detection. To address this issue, we explore two artifact types introduced during the generation process: (1) latent distribution deviations and (2) decoding-induced smoothing effects, which manifest as inconsistencies in local textures, edges, and color transitions. Leveraging local pixel dependencies (LPD) properties rooted in Markov Random Fields, we reconstruct synthetic images using neighboring pixel information to expose disruptions in texture continuity and edge coherence. Building upon LPD, we propose FerretNet, a lightweight neural network with only 1.1M parameters that delivers efficient and robust synthetic image detection. Extensive experiments demonstrate that FerretNet—trained exclusively on the 4-class ProGAN dataset—achieves an average accuracy of 97.1% on an open-world benchmark comprising 22 generative models. Our code and datasets are publicly available at https://github.com/xigua7105/FerretNet.
@@ 1 Introduction
Many existing detection approaches rely heavily on model-specific features, which limit their generalization ability to unseen generative architectures. For example, Durall et al. [8] observed characteristic frequency artifacts in GAN-generated images. Although frequency-domain techniques [51, 16, 17] have demonstrated strong performance under known conditions, they often struggle to generalize across different models. DIRE [52] introduced a diffusion-based detection framework that distinguishes synthetic images by reconstructing them through a diffusion model, a capability that fails with real images. However, this method performs poorly when applied to GAN-generated content.
@@
Based on these insights and grounded in the theory of Markov Random Fields, we introduce a pixel-level artifact representation that captures local pixel dependencies (LPD) through median-based reconstruction. We further propose FerretNet, a lightweight detector incorporating depthwise separable and dilated convolutions to achieve a balance between computational efficiency and representational power.
@@
  • We present Synthetic-Pop, a 60K-image benchmark for evaluating detection models against high-fidelity generators, containing 30K synthetic images from six models and 30K real images from COCO [25] and LAION-Aesthetics V2 (6.5+) [45]. See Appendix A for more details.
@@ 2.1 Pixel-based Synthetic Image Detection
Wang et al. [51] trained a classifier on images generated by a single model to detect fake images across various architectures and datasets, addressing cross-model generalization via data augmentation and diverse training samples. Shi et al. [47] proposed a difference-guided reconstruction learning framework that exploits discrepancies between real and synthetic images to enhance detection accuracy. Ojha et al. [31] tackled the generalization problem to unseen generative models by leveraging a feature space not explicitly trained for real/fake discrimination, employing nearest-neighbor and linear probing strategies. He et al. [13] introduced a super-resolution-based re-synthesis technique to reconstruct test images and extract residual or layered artifact features, thereby reducing reliance on frequency artifacts. Tan et al. [50] proposed NPR, a method that revisits the upsampling process in generative CNNs by modeling Neighbor Pixel Relations, aiming to improve generalization in deepfake detection. Liu et al. [26] designed a robust detection framework based on multi-view image completion, which simulates real image distributions and captures frequency-independent features. FatFormer [27] presented a forgery-aware adaptive transformer incorporating forgery-specific adapters and language-guided alignment modules to better adapt pre-trained models for synthetic image detection. CO-SPY [2] leverages a frozen CLIP encoder for semantic features and a VAE-based reconstruction difference for artifacts, integrating them via adaptive fusion for robust synthetic image detection.
@@ 2.2 Frequency-based Synthetic Image Detection
F3Net [37] introduced a dual-branch architecture that captures frequency-aware clues for detecting subtle forgery traces, particularly in low-quality and facial imagery. FrePGAN [17] developed a frequency-level perturbation GAN framework, where a generator-discriminator pair is used to iteratively improve classifier robustness against unseen categories and generative models. Tan et al. [48] exploited pre-trained CNN gradients to generate generalizable representations of GAN-specific artifacts. BiHPF [16] amplified frequency-level artifacts via a high-pass filtering approach, achieving improved robustness across diverse image categories, color manipulations, and generative models. FreqNet [49] introduced high-frequency representations and frequency-specific convolution layers to enhance detection by focusing on localized high-frequency components, addressing overfitting and poor generalization seen in prior methods. SAFE [24] leverages the high-frequency component of the Discrete Wavelet Transform to extract forensic artifacts and employs data augmentation techniques including ColorJitter, RandomRotation, and a patch-based RandomMask mechanism to improve the model’s generalization and robustness.
@@ 3 Artifacts in Synthetic Image Generation
This section provides a high-level, intuitive framework to motivate our search for universal artifacts. We establish the general principle that all generative models, despite architectural differences, introduce artifacts. While these artifacts have multiple high-level sources (e.g., latent space, decoding), our work chooses to focus on detecting a powerful and universal effect—the disruption of local pixel statistics—which is effectively captured by our LPD method.
@@ 4.1 Local Median-based Feature Extraction
Natural images exhibit strong local statistical consistency due to the underlying physics of light and matter, where neighboring pixels are highly correlated. Generative models, however, often struggle to perfectly replicate these subtle, complex statistics during the synthesis of high-frequency details. This struggle leads to microscopic disruptions in local pixel dependencies. We propose a synthetic image detection method based on local statistical dependencies. The core idea is to identify generation artifacts by quantifying the deviation of each pixel from the median of its surrounding neighborhood. The full computational procedure is outlined in Algorithm 1.
@@ 5.1.1 Training Dataset
To ensure a consistent evaluation baseline, we follow the protocols established in [16, 31, 27, 50], utilizing four semantic classes (car, cat, chair, horse) from the ForenSynths dataset [51]. Each class contains 18,000 synthetic images generated by ProGAN [19], paired with an equal number of real images from the LSUN dataset [56]. All methods compared in this study were trained or fine-tuned on this same limited ProGAN 4-class dataset, except for CO-SPY [2], which utilized its officially released weights trained on other datasets.
@@ 5.3 Main Results
Dataset Wang [51] LGrad [48] Ojha [31] FreqNet [49] NPR [50] FatFormer [27] SAFE [24] CO-SPY [2] FerretNet
Dall-E 51.8/61.3 88.5/97.3 89.5/96.8 97.3/99.7 90.9/98.1 98.8/99.8 97.5/99.7 81.8/87.2 91.4/98.2
Guided 54.9/66.6 86.6/100.0 75.7/85.1 67.2/75.4 74.0/78.1 76.1/92.0 82.4/95.8 62.5/86.0 92.1/98.6
PNDM 50.8/90.3 69.8/98.5 75.3/92.5 85.2/99.9 97.5/100.0 99.3/100.0 78.9/98.6 53.0/55.6 96.9/100.0
VQ-Diffusion 50.0/71.0 96.3/100.0 83.5/97.7 100.0/100.0 100.0/100.0 100.0/100.0 100.0/100.0 71.9/71.5 99.9/100.0
Glide-50-27 54.2/76.0 90.7/95.1 91.1/97.4 86.6/95.8 97.5/99.5 94.7/99.4 96.6/99.2 69.1/74.6 97.2/99.7
Glide-100-10 53.3/72.9 89.4/94.9 90.1/97.0 87.8/96.0 97.8/99.5 94.2/99.2 97.3/99.4 76.6/81.6 97.9/99.9
Glide-100-27 53.0/71.3 87.4/93.2 90.7/97.2 84.4/95.6 97.4/99.5 94.4/99.1 95.8/98.9 73.5/78.2 97.3/99.7
LDM-100 51.9/63.7 94.8/99.2 90.5/97.0 97.8/99.9 98.0/99.6 98.7/99.9 98.8/100.0 82.7/86.9 98.8/100.0
LDM-200 52.0/64.5 94.2/99.1 90.2/97.1 97.4/99.9 98.2/99.6 98.6/99.8 98.8/100.0 83.1/87.5 98.8/100.0
LDM-200-CFG 51.6/63.1 95.9/99.2 77.3/88.6 97.3/99.9 98.0/99.5 94.9/99.1 98.7/99.9 85.3/91.0 98.5/99.9
Mean 52.4/70.1 89.4/97.7 85.4/94.6 90.1/96.2 94.9/97.3 95.0/98.8 94.5/99.1 73.9/80.0 96.9/99.6
@@
Methods Ref Image size Params (M) $\downarrow$ FLOPs (G) $\downarrow$ FPS $\uparrow$ ACC / AP $\uparrow$
FreqNet [49] AAAI 2024 2562 1.85 2.58 200.2 79.2 / 86.8
NPR [50] CVPR 2024 2562 1.44 2.29 720.9 86.5 / 89.4
FatFormer [27] CVPR 2024 2242 492.59 269.92 88.6 86.1 / 91.0
SAFE [24] KDD 2025 2562 1.44 2.29 770.2 96.8 / 99.3
CO-SPY [2] CVPR 2025 3842 963.05 644.80 26.3 76.5 / 83.8
FerretNet (Ours) - 2562 1.06 2.38 772.1 97.1 / 99.6
@@
Next, on diffusion-generated images (Table 2), FerretNet attains an ACC of 96.9% and an AP of 99.6%, outperforming FatFormer [27] by 1.9 and 0.8 percentage points (pp), respectively. Other lightweight models such as NPR [50], FreqNet [49] and SAFE [24] perform less favorably, with ACC scores falling below 95.0%.
We further evaluate performance on high-quality synthetic images using the Synthetic-Pop test set (Table 3). Some existing methods experience noticeable degradation; for example, NPR [50] achieves only 77.4% ACC and 81.2% AP. In contrast, FerretNet maintains 98.3% ACC and 99.8% AP, highlighting its robustness and reliability on visually realistic forgeries.
To evaluate real-world applicability, we tested FerretNet on four distinct test sets for both detection performance and efficiency. As shown in Table 4, FerretNet achieves 97.1% ACC and 99.6% AP with 1.06M parameters and 772.1 FPS on an RTX 4090. Notably, it outperforms FatFormer [27] by 11.0 and 8.6 pp in ACC and AP, respectively, while using only 0.2% of its parameters.
Finally, Appendix B provides a detailed analysis of specific success and failure cases, further clarifying the model’s decision boundaries.
@@ 5.4 Ablation Study
Unless specified, all ablation results report the average ACC and AP across four datasets: ForenSynths, Diffusion-6-cls, Synthetic-Pop, and Synthetic-Aesthetic. More supplementary experiments see Appendix C.
@@ 6 Conclusion
This work presents a universal artifact representation framework and introduces FerretNet, a lightweight yet effective neural network for synthetic image detection. FerretNet achieves a remarkable 99.8% reduction in parameters compared to the state-of-the-art method FatFormer [27], while maintaining exceptional detection accuracy, reaching 97.1% on images generated by 22 different generative models. It demonstrates strong generalization capabilities and computational efficiency, outperforming existing approaches on high-quality synthetic datasets. Our contributions include a novel artifact representation approach and the introduction of the Synthetic-Pop dataset.
@@
Acknowledgements
This work was supported by the National Key R&D Program of China (2022YFC3303603), the National Natural Science Foundation of China (NSFC, 62377028), the Guangdong Basic and Applied Basic Research Foundation (2023B1515120064), the Guangzhou Science and Technology Planning Project (Nansha District: 2023ZD001), the Guangzhou Development District International Cooperation Project (Grant No. 2023GH01), and the Fundamental Research Funds for the Central Universities (21625102).
Appendix
Appendix A Details of Synthetic-Pop
To evaluate the practicality of our method across mainstream generated models, we construct a benchmark named Synthetic-Pop, which includes six widely used models: Openjourney [36], Proteus-0.3 [6], RealVisXL-4.0 [46], SD-3.5-Medium [9], SDXL-Turbo [34], and YiffyMix [55]. Images are generated using 5,000 captions randomly sampled from COCO [25], following the inference configurations recommended by each model. Representative examples are shown in Figure 4.
Openjourney: A bench sitting
on the beach near the ocean.
Proteus-0.3: A brown white
and black dog is laying on a
gray couch.
RealVisXL-4.0: A man that has
glasses and a hat.
SD-3.5-Medium: Two stuffed
animals sit at a table with honey.
SDXL-Turbo: Cat sitting up
with a fake tie around its neck.
YiffyMix: A woman walking
down a street talking on a
cell phone.
Figure 4: Examples of images generated by different models along with their corresponding text prompts. Each subfigure presents an image produced by a specific model, where the format “Model: Prompt” denotes the generating model and its input description.
Appendix B Visualization Analysis
B.1 Success Case Analysis
The effectiveness of LPD in distinguishing synthetic from real images stems from its ability to exploit intrinsic discrepancies in their local statistical structures. First, LPD captures subtle but systematic deviations introduced during the generative process. Although synthetic images may appear perceptually indistinguishable from real images, their micro-texture distributions and noise characteristics deviate from the stochastic sensor noise inherent to real image acquisition. As illustrated in the second and fifth rows of Figure 5, real images retain structured yet naturally irregular noise patterns, whereas synthetic counterparts exhibit overly smooth regions or artificial regularities—a direct consequence of the generative model’s learned priors.
Second, LPD operates in a content-agnostic manner. It consistently reveals statistical inconsistencies across diverse semantic domains, including human portraits, wildlife, and landscapes. This invariance to semantic content highlights the signal-level nature of LPD, ensuring robustness against the rapid evolution of generative models that continue to enhance perceptual fidelity but still leave detectable low-level statistical artifacts.
Finally, the Grad-CAM visualizations in the third and sixth rows confirm that the model’s attention aligns closely with LPD activation regions. Rather than focusing on semantic objects, FerretNet concentrates on areas exhibiting statistical anomalies, enabling consistent detection across diverse image categories. This strong alignment between LPD and Grad-CAM underpins the discriminative strength of our approach: the network remains nearly silent on authentic images while responding sharply to synthetic ones.
Real
LPD
CAM
Fake
LPD
CAM
Figure 5: LPD and Grad-CAM visualizations of real and fake images.
B.2 Failure Case Analysis
False Positive Case. Figure 6 illustrates representative real images are mistakenly classified as synthetic. Unlike the clear statistical distinctions observed in successful detections, these cases demonstrate that certain real-world textures can exhibit statistical patterns similar to those introduced by generative models.
In the first row, geological landscapes display highly regular and repetitive patterns, such as stratified rock formations with sharp color transitions and well-structured edges. These natural textures lead the LPD to capture statistical irregularities that resemble those typically found in generated content, particularly around boundary regions.
In the second row, the image of a bear on grassland demonstrates a similar issue. The network has learned to associate strong, structured statistical biases as primary indicators of artificial generation, resulting in misclassification. This reflects the model’s high sensitivity to subtle statistical cues, while also highlighting its limitations when encountering rare real-world scenes that are themselves highly structured “outliers.”
Figure 6: LPD and Grad-CAM visualizations of False Positive Case.
Figure 7: LPD and Grad-CAM visualizations of False Negative Case.
False Negative Case. Figure 7 illustrates representative synthetic images that were mistakenly classified as real. These cases reveal two distinct mechanisms by which AI-generated content can evade detection, highlighting limitations of LPD-based approaches.
In the synthetic cow image, the LPD map correctly highlights widespread, high-energy artifacts. Here, the failure is not in feature detection but in decision aggregation: the network sums all detected artifact signals to compute a final "fakeness" score. In this instance, despite numerous detected artifacts, their combined contribution did not exceed the threshold for a "fake" classification. This represents a rare boundary-case where a state-of-the-art generator produces content lying precisely on the “real” side of the learned decision boundary.
In contrast, the synthetic giraffe image exhibits a fundamentally different failure mode. Although Grad-CAM shows activation in textured regions, the overall signal is weak and insufficient for confident classification. A major factor is the clean, homogeneous sky background, which occupies a large portion of the image. Such smooth, featureless regions lack the local pattern variations necessary for LPD to detect pixel discontinuities or unnatural transitions. Consequently, the model receives insufficient discriminative signal, leading to misclassification.
Appendix C Additional Experiments
C.1 Scaling FerretNet
We scaled our model down (FerretNet-S) and up (FerretNet-L,  20x parameters). As shown in Table 9, making the model significantly larger results in a slight performance degradation. This confirms our hypothesis that for detecting low-level statistical artifacts, larger networks are prone to overfitting to training-specific patterns, which harms generalization. Our proposed FerretNet hits the "sweet spot" of being sufficiently expressive without the excess capacity.
Table 9: Ablation on scaling FerretNet. Results are mean ACC/AP across all test sets.
Method Channels Blocks Parameters Mean ACC / AP
FerretNet-S (32, 64) (2, 2) 0.13 M 93.1 / 97.6
FerretNet-B (96, 192) (2, 2) 1.06 M 97.1 / 99.6
FerretNet-L (96, 192, 384, 768) (2, 2, 6, 2) 21.51 M 96.6 / 99.4
C.2 Robustness to Common Post-Processing
We conducted a comprehensive robustness analysis on the ForenSynths [51] test set against JPEG compression, resizing, and rotation. For resizing, we used a stringent protocol with dynamic resolutions. As shown in Table 10, FerretNet demonstrates strong robustness, particularly against rotation, where it significantly outperforms the heavyweight FatFormer [27]. This provides strong evidence that LPD features, being based on local, orientation-agnostic statistics, are inherently immune to geometric transformations. While heavy JPEG compression remains a challenge for all lightweight detectors, FerretNet performs competitively.
Table 10: Robustness analysis on the ForenSynths test set against common post-processing attacks.
Method No Attack JPEG Resize Rotation
(Q=100) (Q=75) (S=0.75) (S=1.25) D=[-45°, 45°]
FreqNet [49] 91.5/98.5 50.5/66.6 50.1/51.8 65.2/85.8 64.9/82.8 79.9/91.6
NPR [50] 92.5/96.1 55.0/59.3 50.0/49.1 83.9/84.9 78.9/81.8 86.7/90.7
FatFormer [27] 98.4/99.7 96.5/99.4 71.7/89.8 68.1/96.8
FerretNet (Ours) 95.9/99.3 55.1/67.8 50.2/49.4 81.4/94.3 80.8/95.4 88.2/98.0